OpenStudy (anonymous):

the area of a parallelogram is 32 sq miles. Find its base and height

OpenStudy (anonymous):

the base is set at x+7 and the height is set at x-7

OpenStudy (anonymous):

A=b x h 32=(x+7) (x-7) 32=x^2-7x+7x-49 32=x^2-49 (the 7'scancelled each other out) +49 +49 81=x^2 (sqaure root it) 9=x

jimthompson5910 (jim_thompson5910):

be careful, if you had x^2 = 81, then x = 9 or x = -9

OpenStudy (anonymous):

so then you just plug x into the original equations

jimthompson5910 (jim_thompson5910):

however, in this case, it doesn't make sense to have x = -9 since this leads to a side length of x+7 = -9+7 = -2 so you can get away with ignoring x = -9